OpenStudy (anonymous):

find the derivative of the given function. Simplify and express the answer using positive exponents only. g(x)= (9)/(2(x^3-7)^5)

OpenStudy (anonymous):

make that \[\frac{9}{2}(x^2-7)^{-5}\]

OpenStudy (anonymous):

\[g(x)=\frac{ 9 }{ 2(x ^{3}-7)^{5} }\]

OpenStudy (anonymous):

ok start with \[\frac{9}{2}(x^3-7)^{-5}\]

OpenStudy (anonymous):

clear how to proceed?

OpenStudy (anonymous):

\[\frac{ -45 }{ 2} (x ^{3}-7)^{-6}\]

OpenStudy (anonymous):

then i take the derivative of inside the parentasis?

OpenStudy (anonymous):

@satellite73 \[\frac{ -135 }{ 2 }x ^{2}(x ^{3}-7)^{-6}\] what do i do from here?

OpenStudy (anonymous):

\[-\frac{ 135x ^{2} }{ 2(x ^{3}-7)^{6} }\]

OpenStudy (anonymous):

i guess you can write it back in fraction form

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

yes, that is it! yw (actually you did all the work)

OpenStudy (anonymous):

lol haha yeah but if you hadn't shown me how to re-write the function i would have been clueless! @satellite73 :D